OpenStudy (ksaimouli):

for the circle x^2+y^2+6x-4y+3=0 find the center and radius

OpenStudy (anonymous):

x^2+y^2+6x-4y+3=0 x^2 +6x +9-9 + y^2 -4y -2 +2 +3=0 (x+3)^2 + (y-2)^2 = 9-2-3=4 so center at (-3,2) and radius =2

OpenStudy (ksaimouli):

the equation of the tangent at (-2,5)

OpenStudy (ksaimouli):

how to find this

OpenStudy (anonymous):

sry, made a mistake: x^2+y^2+6x-4y+3=0 x^2 +6x +9-9 + y^2 -4y +4 -4 +3=0 (x+3)^2 + (y-2)^2 = 9+4-3=10 so center at (-3,2) and radius =sqrt10

OpenStudy (anonymous):

\[x^{2}+y^{2}+6x-4y+3\] --> so just move the x's and y's together and put the constant on the other side \[x^{2}+6x+y^{2}-4y=-3\] [ (x^{2}+6x+9)+(y^{2}-4y+4)=-3+4+9\]

OpenStudy (ksaimouli):

i got u @Fgcbear16 and @myko but now i stuck with the tangent

OpenStudy (ksaimouli):

the equation of the tangent at (-2,5)

OpenStudy (anonymous):

good luck!

OpenStudy (anonymous):

x^2+y^2+6x-4y+3=0 derivate this implicitly: 2x+2yy' +6-4y'=0 put terms with y' one side and the rest other: y'(2y-4)=-2x-6 y'=(-2x-6)/(2y-4) substitute value of the point (-2,5) y'=(4-6)/(10-4)=-2/6=-1/3 this is your slope. I think you can finish alone,:)

OpenStudy (ksaimouli):



OpenStudy (anonymous):